# MuonEq: Balancing Before Orthogonalization with Lightweight Equilibration

### 30 Mar 2026 v1

@@ Appendix D Normalization and whitening
  2. (b)
    In the row/left setting of Proposition 3.2, if $\|\mathbf{C}_{r}\|_{2}$ is small enough, then $\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\tfrac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$, so $\|\operatorname{Orth}_{\ell}(\mathbf{N}_{r})-\mathbf{N}_{r}\|_{2}\leq\tfrac{1}{2}\|\mathbf{N}_{r}\|_{2}\|\mathbf{C}_{r}\|_{2}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$. Also,
@@
For (a), Lemma B.1 with $\mathbf{A}=\mathbf{I}_{n}$ and $\mathbf{E}=\mathbf{C}_{c}$ gives the Sylvester equation $\mathbf{L}+\mathbf{L}=\mathbf{C}_{c}$, hence $\mathbf{L}=\tfrac{1}{2}\mathbf{C}_{c}$ and $(\mathbf{I}_{n}+\mathbf{C}_{c})^{-1/2}=\mathbf{I}_{n}-\tfrac{1}{2}\mathbf{C}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$. Left multiplication by $\mathbf{N}_{c}$ gives the expansion of $\operatorname{Orth}_{r}(\mathbf{N}_{c})$, and the norm bound follows from submultiplicativity.
For the comparison with $\mathbf{M}$, Proposition 3.2 gives $\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2})$. Since $\mathbf{M}=\mathbf{N}_{c}\mathbf{D}_{c}$,
@@
Lemma B.2 yields $\|\mathbf{L}_{c}\mathbf{D}_{c}^{-1}\|_{2}\leq\tfrac{1}{2}\kappa(\mathbf{D}_{c})^{2}\|\mathbf{C}_{c}\|_{2}$, which gives the estimate.
The proof of (b) is the same. Lemma B.1 with $\mathbf{A}=\mathbf{I}_{m}$ and $\mathbf{E}=\mathbf{C}_{r}$ gives $(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}=\mathbf{I}_{m}-\tfrac{1}{2}\mathbf{C}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$, hence $\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\tfrac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$. The norm bound again follows from submultiplicativity. Proposition 3.2 also gives $\operatorname{Orth}_{\ell}(\mathbf{M})=\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2})$, so
@@
Using $\|\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\|_{2}\leq\tfrac{1}{2}\kappa(\mathbf{D}_{r})^{2}\|\mathbf{C}_{r}\|_{2}$ completes the proof. ∎
Remark D.1.
Proposition D.1 shows that one-sided normalization removes the zeroth-order marginal scale terms $\|\mathbf{N}_{c}\|_{2}\|\mathbf{D}_{c}-\mathbf{I}_{n}\|_{2}$ and $\|\mathbf{D}_{r}-\mathbf{I}_{m}\|_{2}\|\mathbf{N}_{r}\|_{2}$. Whitening then acts only on the first-order Gram error. Since $\operatorname{diag}(\mathbf{N}_{c}^{\top}\mathbf{N}_{c})=\mathbf{I}_{n}$ and $\operatorname{diag}(\mathbf{N}_{r}\mathbf{N}_{r}^{\top})=\mathbf{I}_{m}$, this leading correction is purely off-diagonal.
Corollary D.2.
@@
  1. (a)
    If $\mathbf{M}$ has full column rank and $\hat{\mathbf{C}}_{c}:=\hat{\mathbf{M}}^{\top}\hat{\mathbf{M}}-\mathbf{I}_{n}$, then for $\|\hat{\mathbf{C}}_{c}\|_{2}$ small enough, $\operatorname{Orth}_{r}(\hat{\mathbf{M}})=\hat{\mathbf{M}}(\mathbf{I}_{n}+\hat{\mathbf{C}}_{c})^{-1/2}=\hat{\mathbf{M}}-\tfrac{1}{2}\hat{\mathbf{M}}\hat{\mathbf{C}}_{c}+\mathcal{O}(\|\hat{\mathbf{C}}_{c}\|_{2}^{2})$, so $\|\operatorname{Orth}_{r}(\hat{\mathbf{M}})-\hat{\mathbf{M}}\|_{2}\leq\tfrac{1}{2}\|\hat{\mathbf{M}}\|_{2}\|\hat{\mathbf{C}}_{c}\|_{2}+\mathcal{O}(\|\hat{\mathbf{C}}_{c}\|_{2}^{2})$.
  2. (b)
    If $\mathbf{M}$ has full row rank and $\hat{\mathbf{C}}_{r}:=\hat{\mathbf{M}}\hat{\mathbf{M}}^{\top}-\mathbf{I}_{m}$, then for $\|\hat{\mathbf{C}}_{r}\|_{2}$ small enough, $\operatorname{Orth}_{\ell}(\hat{\mathbf{M}})=(\mathbf{I}_{m}+\hat{\mathbf{C}}_{r})^{-1/2}\hat{\mathbf{M}}=\hat{\mathbf{M}}-\tfrac{1}{2}\hat{\mathbf{C}}_{r}\hat{\mathbf{M}}+\mathcal{O}(\|\hat{\mathbf{C}}_{r}\|_{2}^{2})$, so $\|\operatorname{Orth}_{\ell}(\hat{\mathbf{M}})-\hat{\mathbf{M}}\|_{2}\leq\tfrac{1}{2}\|\hat{\mathbf{M}}\|_{2}\|\hat{\mathbf{C}}_{r}\|_{2}+\mathcal{O}(\|\hat{\mathbf{C}}_{r}\|_{2}^{2})$.
@@
Appendix E Lemmas for Proposition 3.4
E.1 Lemma E.1
Lemma E.1.
Let the momentum iterate satisfy $\mathbf{M}_{t}=\beta\mathbf{M}_{t-1}+(1-\beta)\mathbf{G}_{t},\qquad\beta\in[0,1)$, and assume there exists a constant $G_{\infty}>0$ such that $\|\mathbf{G}_{t}\|_{\infty}\leq G_{\infty},\|\mathbf{M}_{0}\|_{\infty}\leq G_{\infty}$. For $\varepsilon>0$, define $\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})+\varepsilon\bigr),\mathbf{D}_{c,t}:=\operatorname{diag}\bigl(\operatorname{colsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})+\varepsilon\bigr)$, and $\mathbf{P}_{t}:=\mathbf{D}_{r,t}^{-1/2},\mathbf{Q}_{t}:=\mathbf{D}_{c,t}^{-1/2}$. Then, for all $t$,
@@
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}a^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$
@@
$+\beta_{t+1}^{2}\left(1+\frac{1}{\epsilon}\right)L^{2}\eta_{t}^{2}a^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$
@@
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}a^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$ (6)
Appendix F Proofs of Proposition 3.4
@@
First, by the singular-value inequality
@@
Second, using Hölder’s inequality, $\|\mathbf{O}_{t}\|_{2}=1$, and the bounds above,
@@
Third, we use the symmetric decomposition
@@
According to Lemma E.2, by letting $E_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},A_{t+1}=h_{t+1}^{2}(2\sqrt{2}L^{2}a^{2}n+\sigma^{2})$, we have
@@
Appendix G Lemmas for Theorem 3.5
G.1 Lemma G.1
Lemma G.1.
Let $\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})\bigr),\mathbf{R}_{t}:=\mathbf{D}_{r,t}^{-1/2}\mathbf{M}_{t},\mathbf{O}_{t}:=\operatorname{Orth}(\mathbf{R}_{t})$. Then, for all $t$,
@@
If $\mathbf{M}_{t}=\mathbf{0}$, then $\mathbf{R}_{t}=\mathbf{O}_{t}=\mathbf{0}$ and the claim is immediate. Assume $\mathbf{M}_{t}\neq\mathbf{0}$. Write
@@
G.2 Lemma G.2
Consider the practical row-normalized R variant (4), with Nesterov disabled and $\lambda=0,\varepsilon=0$. Then the update is
$\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})\bigr),\hat{\mathbf{M}}_{t}:=(\mathbf{D}_{r,t}^{1/2})^{\dagger}\mathbf{M}_{t},\mathbf{O}_{t}:=\mathrm{NS5}(\hat{\mathbf{M}}_{t}),\mathbf{X}_{t+1}=\mathbf{X}_{t}-a\eta_{t}\mathbf{O}_{t},$
where $a:=0.2\sqrt{\max\{m,n\}}$. Here denotes the Moore–Penrose pseudoinverse, so a zero row remains zero after row normalization, and we adopt the convention $\operatorname{Orth}(\mathbf{0})=\mathbf{0}$.
Remark G.1.
Following the practical Muon implementation [20] and Theorem 3.1, we write $\mathrm{NS5}$ as the five-step pre-scaled Newton–Schulz map. For any $\mathbf{A}\in\mathbb{R}^{m\times n}$, define $\alpha(\mathbf{A}):=\max\{1,\|\mathbf{A}\|_{F}\},\mathbf{Y}^{(0)}(\mathbf{A}):=\mathbf{A}/\alpha(\mathbf{A}),$ and, for $k=0,1,2,3,4$,
$\mathbf{Y}^{(k+1)}(\mathbf{A})=p_{\mathrm{ns}}\bigl(\mathbf{Y}^{(k)}(\mathbf{A})\mathbf{Y}^{(k)}(\mathbf{A})^{\top}\bigr)\mathbf{Y}^{(k)}(\mathbf{A}),\qquad p_{\mathrm{ns}}(\mathbf{Z}):=a_{\mathrm{ns}}\mathbf{I}_{m}+b_{\mathrm{ns}}\mathbf{Z}+c_{\mathrm{ns}}\mathbf{Z}^{2}.$
We then define
@@
Remark G.2.
Following the trajectory-wise error formalism in [17, Definition 3 and Theorem 2], we work with the NS5 polar-approximation error along the algorithmic trajectory. We define
$\varepsilon_{\mathrm{ns}}:=\sup_{t\geq 1}\bigl\|\mathrm{NS5}(\hat{\mathbf{M}}_{t})-\operatorname{Orth}(\hat{\mathbf{M}}_{t})\bigr\|_{2}.$
In subsequent derivations, we assume $\varepsilon_{\mathrm{ns}}<1$. The quantity $\varepsilon_{\mathrm{ns}}$ plays the same role as the trajectory-wise polar-approximation error $\varepsilon_{q}$ in [17, Definition 3 and Theorem 2].
Lemma G.2.
@@
Consequently, every diagonal entry $\tilde{\sigma}_{i}$ of $\widetilde{\boldsymbol{\Sigma}}$ satisfies
$1-\varepsilon_{\mathrm{ns}}\leq\tilde{\sigma}_{i}\leq 1+\varepsilon_{\mathrm{ns}},\qquad i\in[\operatorname{rank}(\mathbf{A})].$
@@
$\max_{i}|\tilde{\sigma}_{i}-1|\leq\varepsilon_{\mathrm{ns}},$
@@
$1-\varepsilon_{\mathrm{ns}}\leq\tilde{\sigma}_{i}\leq 1+\varepsilon_{\mathrm{ns}}\qquad\text{for all }i.$
@@
G.3 Lemma G.3
Lemma G.3.
Under Lemma G.2, for all $t$,
@@
For the norm bound, let $r_{t}:=\operatorname{rank}(\hat{\mathbf{M}}_{t})$. Again by Lemma G.2, every nonzero singular value of $\mathbf{O}_{t}$ is at most $1+\varepsilon_{\mathrm{ns}}$. Therefore
$\|\mathbf{O}_{t}\|_{F}^{2}=\sum_{i=1}^{r_{t}}\tilde{\sigma}_{i,t}^{2}\leq(1+\varepsilon_{\mathrm{ns}})^{2}r_{t}\leq(1+\varepsilon_{\mathrm{ns}})^{2}n.$
@@
Appendix H Proofs of Theorem 3.5

### 10 May 2026 v2

@@ Appendix E Normalization and whitening
In the row/left setting of Proposition 3.2, if $\|\mathbf{C}_{r}\|_{2}$ is sufficiently small, then $\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\tfrac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$, and therefore $\|\operatorname{Orth}_{\ell}(\mathbf{N}_{r})-\mathbf{N}_{r}\|_{2}\leq\tfrac{1}{2}\|\mathbf{N}_{r}\|_{2}\|\mathbf{C}_{r}\|_{2}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$. Moreover,
@@
Consider first the column/right case. Applying Lemma C.1 with $\mathbf{A}=\mathbf{I}_{n}$ and $\mathbf{E}=\mathbf{C}_{c}$ gives the Sylvester equation $\mathbf{L}+\mathbf{L}=\mathbf{C}_{c}$. Thus $\mathbf{L}=\tfrac{1}{2}\mathbf{C}_{c}$, and $(\mathbf{I}_{n}+\mathbf{C}_{c})^{-1/2}=\mathbf{I}_{n}-\tfrac{1}{2}\mathbf{C}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$. Multiplying on the left by $\mathbf{N}_{c}$ gives the expansion of $\operatorname{Orth}_{r}(\mathbf{N}_{c})$, and the norm bound follows from submultiplicativity.
To compare with $\mathbf{M}$, Proposition 3.2 gives $\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2})$. Since $\mathbf{M}=\mathbf{N}_{c}\mathbf{D}_{c}$,
@@
Lemma C.2 yields $\|\mathbf{L}_{c}\mathbf{D}_{c}^{-1}\|_{2}\leq\tfrac{1}{2}\kappa(\mathbf{D}_{c})^{2}\|\mathbf{C}_{c}\|_{2}$, which gives the stated estimate.
The row/left case follows by the same argument. Applying Lemma C.1 with $\mathbf{A}=\mathbf{I}_{m}$ and $\mathbf{E}=\mathbf{C}_{r}$ gives $(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}=\mathbf{I}_{m}-\tfrac{1}{2}\mathbf{C}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$, and hence $\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\tfrac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$. The norm bound again follows from submultiplicativity. Proposition 3.2 also gives $\operatorname{Orth}_{\ell}(\mathbf{M})=\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2})$, so
@@
The bound $\|\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\|_{2}\leq\tfrac{1}{2}\kappa(\mathbf{D}_{r})^{2}\|\mathbf{C}_{r}\|_{2}$ completes the proof. ∎
Remark E.1.
Proposition E.1 shows that one-sided normalization removes the zeroth-order marginal scale terms $\|\mathbf{N}_{c}\|_{2}\|\mathbf{D}_{c}-\mathbf{I}_{n}\|_{2}$ and $\|\mathbf{D}_{r}-\mathbf{I}_{m}\|_{2}\|\mathbf{N}_{r}\|_{2}$. Whitening then acts only on the first-order Gram error. Since $\operatorname{diag}(\mathbf{N}_{c}^{\top}\mathbf{N}_{c})=\mathbf{I}_{n}$ and $\operatorname{diag}(\mathbf{N}_{r}\mathbf{N}_{r}^{\top})=\mathbf{I}_{m}$, the leading correction is purely off-diagonal.
Corollary E.2.
@@
If $\mathbf{M}$ has full column rank and $\hat{\mathbf{C}}_{c}:=\hat{\mathbf{M}}^{\top}\hat{\mathbf{M}}-\mathbf{I}_{n}$, then for $\|\hat{\mathbf{C}}_{c}\|_{2}$ sufficiently small, $\operatorname{Orth}_{r}(\hat{\mathbf{M}})=\hat{\mathbf{M}}(\mathbf{I}_{n}+\hat{\mathbf{C}}_{c})^{-1/2}=\hat{\mathbf{M}}-\tfrac{1}{2}\hat{\mathbf{M}}\hat{\mathbf{C}}_{c}+\mathcal{O}(\|\hat{\mathbf{C}}_{c}\|_{2}^{2})$, and therefore $\|\operatorname{Orth}_{r}(\hat{\mathbf{M}})-\hat{\mathbf{M}}\|_{2}\leq\tfrac{1}{2}\|\hat{\mathbf{M}}\|_{2}\|\hat{\mathbf{C}}_{c}\|_{2}+\mathcal{O}(\|\hat{\mathbf{C}}_{c}\|_{2}^{2})$.
If $\mathbf{M}$ has full row rank and $\hat{\mathbf{C}}_{r}:=\hat{\mathbf{M}}\hat{\mathbf{M}}^{\top}-\mathbf{I}_{m}$, then for $\|\hat{\mathbf{C}}_{r}\|_{2}$ sufficiently small, $\operatorname{Orth}_{\ell}(\hat{\mathbf{M}})=(\mathbf{I}_{m}+\hat{\mathbf{C}}_{r})^{-1/2}\hat{\mathbf{M}}=\hat{\mathbf{M}}-\tfrac{1}{2}\hat{\mathbf{C}}_{r}\hat{\mathbf{M}}+\mathcal{O}(\|\hat{\mathbf{C}}_{r}\|_{2}^{2})$, and therefore $\|\operatorname{Orth}_{\ell}(\hat{\mathbf{M}})-\hat{\mathbf{M}}\|_{2}\leq\tfrac{1}{2}\|\hat{\mathbf{M}}\|_{2}\|\hat{\mathbf{C}}_{r}\|_{2}+\mathcal{O}(\|\hat{\mathbf{C}}_{r}\|_{2}^{2})$.
@@
Appendix F Lemmas for Proposition 3.4
F.1 Lemma F.1
Lemma F.1.
Let the momentum iterate satisfy $\mathbf{M}_{t}=\beta\mathbf{M}_{t-1}+(1-\beta)\mathbf{G}_{t},\beta\in[0,1)$, and assume there exists a constant $G_{\infty}>0$ such that $\|\mathbf{G}_{t}\|_{\infty}\leq G_{\infty},\|\mathbf{M}_{0}\|_{\infty}\leq G_{\infty}$. For $\varepsilon>0$, define $\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})+\varepsilon\bigr),\mathbf{D}_{c,t}:=\operatorname{diag}\bigl(\operatorname{colsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})+\varepsilon\bigr)$, and $\mathbf{P}_{t}:=\mathbf{D}_{r,t}^{-1/2},\mathbf{Q}_{t}:=\mathbf{D}_{c,t}^{-1/2}$. Then, for all $t$,
@@
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq$ $\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]$
$+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}^{2}+(1-\beta_{t+1})^{2}\sigma^{2}.$
@@
$+\beta_{t+1}^{2}\left(1+\frac{1}{\epsilon}\right)L^{2}\eta_{t}^{2}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}^{2}+(1-\beta_{t+1})^{2}\sigma^{2}.$
@@
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq$ $\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]$ (6)
$+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}^{2}+(1-\beta_{t+1})^{2}\sigma^{2}.$
Appendix G Proofs of Proposition 3.4
@@
(i) First, by the singular-value inequality
@@
(ii) Second, using Hölder’s inequality, $\|\mathbf{O}_{t}\|_{2}=1$, and the bounds above,
@@
(iii) Third, we use the symmetric decomposition
@@
According to Lemma F.2, by letting $E_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},A_{t+1}=h_{t+1}^{2}(2\sqrt{2}L^{2}a^{2}n+\sigma^{2})$, we have
@@
Appendix H Lemmas for Theorem 3.5 and Corollary 3.6
H.1 Lemma H.1
Lemma H.1.
Let
$\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})\bigr),\quad\mathbf{P}_{r,t}:=(\mathbf{D}_{r,t}^{1/2})^{\dagger},\quad\mathbf{R}_{t}:=\mathbf{P}_{r,t}\mathbf{M}_{t},\quad\mathbf{O}_{t}:=\operatorname{Orth}(\mathbf{R}_{t}),$
where denotes the Moore–Penrose pseudoinverse, so zero rows of $\mathbf{M}_{t}$ remain zero after row normalization, and $\operatorname{Orth}(\mathbf{0}):=\mathbf{0}$. Then, for all $t$,
@@
(i) If $\mathbf{M}_{t}=\mathbf{0}$, then $\mathbf{R}_{t}=\mathbf{O}_{t}=\mathbf{0}$ and the claim is immediate.
(ii) Assume $\mathbf{M}_{t}\neq\mathbf{0}$. Let
@@
H.2 Lemma H.2
Lemma H.2.
Let $\{\hat{\mathbf{M}}_{t}\}_{t=1}^{T}\subset\mathbb{R}^{m\times n}$ be any finite algorithmic trajectory, and set $\operatorname{Orth}(\mathbf{0}):=\mathbf{0}$. Consider the five-step pre-scaled Newton–Schulz map with the degree-two Taylor coefficients analyzed in [18, Definition 2]. For any $\mathbf{A}\in\mathbb{R}^{m\times n}$, define
$\alpha(\mathbf{A}):=\max\{1,\|\mathbf{A}\|_{F}\},\qquad\mathbf{Y}^{(0)}(\mathbf{A}):=\frac{\mathbf{A}}{\alpha(\mathbf{A})},$
and, for $k=0,1,2,3,4$,
$\mathbf{Y}^{(k+1)}(\mathbf{A})=p_{\mathrm{ns}}\bigl(\mathbf{Y}^{(k)}(\mathbf{A})\mathbf{Y}^{(k)}(\mathbf{A})^{\top}\bigr)\mathbf{Y}^{(k)}(\mathbf{A}),$
where
$p_{\mathrm{ns}}(\mathbf{Z}):=\frac{15}{8}\mathbf{I}_{m}-\frac{5}{4}\mathbf{Z}+\frac{3}{8}\mathbf{Z}^{2}.$
We define
@@
For each $t$, let $\Pi_{t}$ be the orthogonal projector onto $\operatorname{range}(\hat{\mathbf{M}}_{t})$, and define
$\delta_{t,0}:=\bigl\|\Pi_{t}-\mathbf{Y}^{(0)}(\hat{\mathbf{M}}_{t})\mathbf{Y}^{(0)}(\hat{\mathbf{M}}_{t})^{\top}\bigr\|_{2},\qquad\delta_{0,T}:=\max_{1\leq t\leq T}\delta_{t,0}.$
Then $\delta_{0,T}<1$. Moreover, the trajectory-wise NS5 polar-approximation error
$\varepsilon_{\mathrm{ns},T}:=\max_{1\leq t\leq T}\bigl\|\mathrm{NS5}(\hat{\mathbf{M}}_{t})-\operatorname{Orth}(\hat{\mathbf{M}}_{t})\bigr\|_{2}<1.$
Proof.
If $\hat{\mathbf{M}}_{t}=\mathbf{0}$, then $\Pi_{t}=\mathbf{0}$ and $\delta_{t,0}=0$. Otherwise, the pre-scaling by $\alpha(\hat{\mathbf{M}}_{t})=\max\{1,\|\hat{\mathbf{M}}_{t}\|_{F}\}$ gives
$0<\lambda_{\min}^{+}\bigl(\mathbf{Y}^{(0)}(\hat{\mathbf{M}}_{t})\mathbf{Y}^{(0)}(\hat{\mathbf{M}}_{t})^{\top}\bigr)\leq 1$
on $\operatorname{range}(\hat{\mathbf{M}}_{t})$, and hence $\delta_{t,0}<1$. Since the horizon $T$ is finite, $\delta_{0,T}=\max_{1\leq t\leq T}\delta_{t,0}<1$.
The polynomial above is the degree-two Taylor Newton–Schulz polynomial in [18, Definition 2]. Applying [18, Theorem 2] with $\kappa=2$ and $q=5$ yields
$\varepsilon_{\mathrm{ns},T}\leq 1-\sqrt{1-\delta_{0,T}^{(\kappa+1)^{q}}}=1-\sqrt{1-\delta_{0,T}^{3^{5}}}.$
Finally, since $1-\sqrt{1-x}\leq x$ for $x\in[0,1]$ and $\delta_{0,T}<1$, we obtain
$\varepsilon_{\mathrm{ns},T}\leq\delta_{0,T}^{3^{5}}<1.$
H.3 Lemma H.3
Lemma H.3.
@@
Consequently, every diagonal entry $\widetilde{\sigma}_{i}$ of $\widetilde{\boldsymbol{\Sigma}}$ satisfies
$1-\varepsilon_{\mathrm{ns}}\leq\widetilde{\sigma}_{i}\leq 1+\varepsilon_{\mathrm{ns}},\qquad i\in[\operatorname{rank}(\mathbf{A})].$
@@
$\max_{i}|\widetilde{\sigma}_{i}-1|\leq\varepsilon_{\mathrm{ns}},$
@@
$1-\varepsilon_{\mathrm{ns}}\leq\widetilde{\sigma}_{i}\leq 1+\varepsilon_{\mathrm{ns}}\qquad\text{for all }i.$
@@
H.4 Lemma H.4
Lemma H.4.
Under Lemma H.3, for all $t$,
@@
For the norm bound, let $r_{t}:=\operatorname{rank}(\hat{\mathbf{M}}_{t})$. Again by Lemma H.3, every nonzero singular value of $\mathbf{O}_{t}$ is at most $1+\varepsilon_{\mathrm{ns}}$. Therefore
$\|\mathbf{O}_{t}\|_{F}^{2}=\sum_{i=1}^{r_{t}}\widetilde{\sigma}_{i,t}^{2}\leq(1+\varepsilon_{\mathrm{ns}})^{2}r_{t}\leq(1+\varepsilon_{\mathrm{ns}})^{2}n.$
@@
H.5 Lemma H.5
Lemma H.5.
Consider the row-normalized update with decoupled weight decay
$\mathbf{X}_{t+1}=\mathbf{X}_{t}-\eta_{t}(a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}),$
where $\mathbf{O}_{t}=\mathrm{NS5}(\hat{\mathbf{M}}_{t})$ and Lemma H.4 holds. Let $\gamma:=1-\varepsilon_{\mathrm{ns}},0\leq\rho<\frac{a\gamma}{\sqrt{m}},$ and assume
$0\leq\lambda_{t}\leq\frac{\rho}{\|\mathbf{X}_{1}\|_{F}+4a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}}t^{-1/4}.$ (7)
Then, along every realization,
$\lambda_{t}\|\mathbf{X}_{t}\|_{F}\leq\rho,\qquad\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}\leq(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho)\cdot\eta_{t}.$